OpenStudy (anonymous):

You drop a rock down a well that is 31 m deep. How long does it take the rock to hit the bottom of the well?

OpenStudy (anonymous):

t = d/v d = 31 v = 9.8

OpenStudy (anonymous):

If the rock is dropped from rest, \[\color{green}{t = \sqrt{\frac{2 \Delta \vec{d}}{\vec{g}}}}\]

OpenStudy (anonymous):

bennyboy489 is assuming that there is no acceleration, so his solution is incorrect.

OpenStudy (anonymous):

ahhhh, ok, was under the assumption it was already falling